# Engineering Carbon Credits Towards A Responsible FinTech Era: The Practices, Implications, and Future

### v1

@@ 7 Prospects of Carbon Credits in the Financial and FinTech Fields
The existing literature has extensively studied the impact of carbon emissions and carbon information disclosure on corporate market value, financial performance, and loan interest rates. However, the role of carbon credits in corporate valuation, particularly in the financial and FinTech fields, remains an area with significant potential for further exploration.
@@ 8 Literature Review Comparison
Currently, there is a very limited number of literature reviews on the topic of carbon credits. Below is a list of the most recent reviews concerning carbon credits and carbon management.
Trouwloon reviewed the use of carbon credits in corporate climate statements, focusing on the classification of corporate climate statements, quality issues related to carbon credits, and the aspects of risk and governance within these statements. By examining the role of carbon credits in corporate climate statements, this study provides a deeper understanding of corporate climate declarations, emphasizing the responsibilities and challenges organisations face in climate action, as well as how transparency and sustainability can be improved [13].
Lokuge and Anders conducted a comprehensive literature review on the carbon credit system in Canada’s agricultural sector, with a particular focus on Alberta. They discussed the impact of carbon tax policies on various industries, the research on global carbon credit markets, and the carbon credit systems in the agricultural sectors of developed countries, with special emphasis on Alberta and the western regions of Canada [14].
Alshatri and Hussain conducted a systematic literature review identifying the external factors influencing carbon credit prices and categorized them into energy market factors, environmental variables, macroeconomic indicators, social factors, and online search indices. They emphasized the complexity and multidimensionality of these factors [2].
Jin et al. systematically reviewed 147 carbon emission prediction models (CEPM) conducted between 2011 and August 2022, with a focus on model classification, optimization techniques, and influencing factors. However, this review primarily focuses on the prediction of carbon emissions at the national, industrial, or city level, rather than at the corporate level, thus providing limited insight for investors and governments to understand the specifics of individual organisations [118].
Wang et al. reviewed the impact of carbon risk on corporate capital costs and market value, highlighting the negative effects of high carbon emissions on corporate financial performance. However, Wang et al.’s review mainly concentrated on the direct impact of carbon emissions on organisations and did not fully explore the role of carbon emission disclosure in corporate management and market response [119].
@@ 9 Conclusion
Finally, integrating carbon price predictions with carbon emission forecasts allows organisations to manage their carbon emission costs more comprehensively. The combination of these two techniques lays the foundation for future quantitative research, providing more precise tools for corporate valuation and market performance, helping organisations achieve sustainable development goals while enhancing market competitiveness.
In this way, these advancements not only assist organisations in optimizing their carbon management strategies but also promote the development of the global carbon credit market, contributing to the achievement of sustainable development goals.

### v2

@@ 7 Prospects of Carbon Credits in the Financial and FinTech Fields
The existing literature has extensively studied the impact of carbon emissions and carbon information disclosure on corporate market value, financial performance, and loan interest rates. However, the role of carbon credits in corporate valuation, particularly in the financial and FinTech fields, remains an under explored area.
@@ 8 Literature Review Comparison
Currently, a comprehensive literature review on the topic of carbon credits is still lacking. Below we summarise a list of the most recent reviews concerning carbon credits and carbon management.
Trouwloon reviewed the use of carbon credits in corporate climate statements, focusing on the classification of corporate climate statements, quality issues related to carbon credits, and the aspects of risk and governance within these statements. By examining the role of carbon credits in corporate climate statements, this study provides a deeper understanding of corporate climate declarations, emphasizing the responsibilities and challenges organisations face in climate action, as well as how transparency and sustainability can be improved [carbon75].
Lokuge and Anders conducted a comprehensive literature review on the carbon credit system in Canada’s agricultural sector, with a particular focus on Alberta. They discussed the impact of carbon tax policies on various industries, the research on global carbon credit markets, and the carbon credit systems in the agricultural sectors of developed countries, with special emphasis on Alberta and the western regions of Canada [carbon76].
Alshatri and Hussain conducted a systematic literature review identifying the external factors influencing carbon credit prices and categorized them into energy market factors, environmental variables, macroeconomic indicators, social factors, and online search indices. They emphasized the complexity and multidimensionality of these factors [carbon2].
Jin et al. systematically reviewed 147 carbon emission prediction models between 2011 and August 2022, with a focus on model classification, optimization techniques, and influencing factors. However, this review primarily focuses on the prediction of carbon emissions at the national, industrial, or city level, rather than at the corporate level, thus providing limited insight for investors and governments to understand the specifics of individual organisations [lr1].
Wang et al. reviewed the impact of carbon risk on corporate capital costs and market value, highlighting the negative effects of high carbon emissions on corporate financial performance. However, Wang et al.’s review mainly concentrated on the direct impact of carbon emissions on organisations and did not fully explore the role of carbon emission disclosure in corporate management and market response [lr2].
@@ 9 Conclusion
Finally, integrating carbon price predictions with carbon emission forecasts allows organisations to manage their carbon emission costs more comprehensively. The combination of these two techniques lays the foundation for future quantitative research, providing more precise tools for corporate valuation and market performance, helping organisations achieve sustainable development goals while enhancing market competitiveness. In this way, these advancements not only assist organisations in optimizing their carbon management strategies but also promote the development of the global carbon credit market, contributing to the achievement of sustainable development goals.